Meeting Notes for 22 May

Present: Bob, Phong, Ken, Liz and Edit


  • New Drupal site: we need to do some research on how to set this up at Phong will check.
  • Stress test on 13 scheduled for today. We will be testing Mihael's changes.
  • Follow-up about disks:
    • data2: Edit will compress LIGO old frames using xz.
  • Bug with pre-tests and post-tests. Needs to be fixed.
-- Main.EditPeronja - 2013-05-29
Topic revision: r1 - 2013-05-29, eperonja
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback